# Constitutional Consensus

### v4

@@ 4.2 Liveness
{observation}
[Single-epoch Liveness] A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
Theorem 4.10.
@@
Proof 4.12.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $b_{p}]$) have not been received by $q$, then after receiving $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction.
@@
{observation}
[Sequential Advance] If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
@@
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation 4.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition 4.11, this means that no correct agent sends blocks in rounds >r>r.
Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition 4.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
By Proposition 4.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
@@
Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction.
@@
Proposition 4.15 (Finite implies quiescence).
@@
Proposition 4.17 (Block issuance).
@@
Proof 4.18.
Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload. First, consider the case that $p$’s rounds do not advance in some suffix of the run. Then by Proposition 4.15, $p$’s blocklace is quiescent. Hence, condition 1 of the Issue rule is satisfied, and the payload is sent. A contradiction. Therefore, $p$’s advanced rounds increase indefinitely. Consider a time when $p$’s highest advanced round $r>r_{0}$ and the payload has not been sent. Then $p$ issues a block either in round $r+1$ (by the Issue rule) or in round $r$ (by the Issue backlog rule). A contradiction.
Proposition 4.19 (Transaction Liveness in finite runs).
@@
Proof 4.20.
By Proposition 4.11, $B$ includes all ordinary blocks sent by correct agents. By Proposition 4.15, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$. Moreover, by Proposition 4.17, no agents have payloads to send. It follows that all input transactions have been sent in ordinary blocks and ordered.
Proposition 4.21 (Epoch Liveness in finite runs).
@@
Proof 4.22.
Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run. Because $t$ is at some point final in $p_{1}$’s blocklace, its blocklace eventually includes a block $b$ with $t$ in its payload and a super-majority of blocks that observe $b$. This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$. By Proposition 4.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$. Because $B_{2}$ is finite in a suffix of the run, by Proposition 4.15, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
@@
Proposition 4.23 (First round progress after GST).
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt>GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof 4.24.
@@
When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block.
@@
Proposition 4.25 (Progress after GST).
@@
Proof 4.26.
By Proposition 4.23, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$. This means that it is impossible to create a valid round $r+1$ block that does not endorse the leader, as such a block would have to point to a super-majority of round $r$ blocks excluding the leader block. Therefore, every valid round $r+2$ block ratifies the leader block, and the block is finalized in round $r+2$ as soon as it advances at any correct agent.
@@
Proposition 4.27 (Infinite quiescent waves).
@@
Proof 4.28.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition 4.11, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalising, and finalization is monotonic, there are infinitely many finalising waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
@@
Consider a finalising wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 4.11). If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
@@
So when the third round of wave $k+1$ advances, every correct agent either observes a super-majority of blocks that ratify the leader block or observes a block whose closure is quiescent. In the former case the leader block is finalized, ordering all preceding ordinary blocks including $b_{2}$, and in the latter $b_{2}$ is already final in the block’s closure.
@@
By Observation 4.2, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent. By Proposition 4.17, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks. Similarly, every output transaction of a correct agent is included in some ordinary block. Let $b$ be a non-empty ordinary $p$-block included in some correct agent’s blocklace in a run of Constitutional Consensus. If $p$ is correct, then by Proposition 4.11, every correct agent receives and incorporates $b$ in its local blocklace. Otherwise, $b$ is finalized in some blocklace, implying that it is approved by at least one $q$-block $b^{\prime}$ by a correct agent $q$. Again, by Proposition 4.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
@@
We consider three cases: First, assume the run is finite. By Propositions 4.19 and 4.21, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds. Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. By Proposition 4.27, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
@@
Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time. Let $k$ be a wave whose last round advances after GST and after which no wave is quiescent at any correct agent’s blocklace. By fairness of leader selection, there are inifinitely many waves $k^{\prime}>k$ with a correct leader. By Proposition 4.25, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
5 Constitutional Consensus with Constitutional Amendments
The problem of reconfiguration of a consensus protocol—namely changing the set of its participants—has engaged the distributed computing research community for decades [40, 42, 2, 40], ever since it was first formulated in Lamport’s Paxos paper [21]. Here we explore the more general problem of constitutional amendments, explained and defined below. Supporting constitutional amendments makes the consensus protocol grassroots, as proven in Section 6.
5.1 Constitutional Consensus: From General to Specific
In its most general sense, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself. The notion of constitutional amendment is different from the well-studied notion of reconfiguration in at least two respects:
  1. 1.
    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including the agents executing it, but also the protocol itself and its parameters.
  2. 2.
    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to descend on obedient agents “out of the blue”. Amending the constitution of a protocol, on the other hand, is done by the volitional agents (e.g., people) executing the protocol, in accordance with the prevailing constitutional amendment rule.
In this work we consider a restricted type of constitution $(P,\sigma,\Delta)$ (Definition 2.1), the components of which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony. Even more so, they are key to Constitutional Consensus, given that its primary aim is to execute digital social contracts [9, 38]:
  1. 1.
    $P$: The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc. Clearly, they should be sovereign to accept new members and remove existing members, namely to amend $P$.
  2. 2.
    $\sigma$: As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should the $\sigma$-supermajority among $P$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased. Adding new people to $P$, initially known to and trusted by only a fraction of $P$, may entail increasing $\sigma$, even if temporarily, in order to ensure that sufficiently many trusted agents are part of every decision.
  3. 3.
    $\Delta$: The delay parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members with lower-capacity devices are admitted, $\Delta$ may need to be increased. In addition, technology advances may allow decreasing $\Delta$. Thus, the prevailing population $P$ should be sovereign to amend the value $\Delta$ as they see fit.
5.2 Democratic Constitutional Amendments
The Constitutional Consensus protocol attempts to implement valid constitutional amendment decisions as soon as they are made, but is independent of how these decisions are made. Computational Social Choice research [6] has investigated democratic decision-making processes extensively. Here we describe possible democratic processes for amending each of the components of a constitution $(P,\sigma,\Delta)$.
(1) Amending $P$. For amending $P$ to $P^{\prime}$, we require that a $\sigma$-supermajority among $P$ and that all new members of $P^{\prime}$, namely those in $P^{\prime}\setminus P$, approve the decision. The purpose of the supermajority is to ensure that a majority of the correct agents in $P$ are willing to hand-over control to $P^{\prime}$. The purpose of requiring all new participants in $P^{\prime}$ to sign is, first, to respect their autonomy—an agent should agree to participate—and, second, to avoid the ‘dead sailors problem’ [21], in which all of $P^{\prime}$ are new but in fact dead, resulting in a constitutional stalemate.
The respect for autonomy of new participants has a practical corollary: An honest person (who would participate as a correct agent) would not join a protocol instance if they are concerned that the prevailing population may have too many Byzantines, namely that the prevailing constitution is not safe. The flip side of this concern—that the prevailing population not add malicious agents—has been studied extensively, as discussed next.
A constitutional consensus protocol shares a fundamental challenge with permissionless consensus protocols: Both allow the participating agent set to change dynamically, yet must prevent Sybil attacks where a single entity creates multiple digital agents to gain disproportionate influence. Permissionless protocols address this by requiring agents to prove control of scarce resources—computational power in Proof-of-Work [27] or financial stake in Proof-of-Stake [19, 8]. We expect realization of constitutional consensus to take a people’s-based approach: To rely on trust relations among real people, and admit new agents only when existing members know, trust, and verify them as genuine individuals. This requires (i) sybil-resistant admission protocols [28, 33] during community formation to ensure that digital agents genuinely represent distinct people rather than fake or duplicate identities and (ii) sybil-resilient democratic governance [32, 24], to ensure that sybils cannot overpower the genuine agents in democratic decisions. These mechanisms are complementary to this work.
(2) Amending $\sigma$. Amending the supermajority by which a constitution can be amended has been studied by Abramowitz et al. [1], who showed that the $h$-rule (similar to the $h$-index of author citation metrics) is the only rule that satisfies certain intuitive and self-evident axioms. According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$, in a process left unspecified. The value of $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which there is a $\sigma^{\prime}$-supermajority among the prevailing population $P$ who voted for a value greater or equal to $\sigma^{\prime}$. Else $\sigma$ is decreased to the minimal $\sigma^{\prime}$ for which there is a $\sigma$-supermajority among the prevailing population $P$ who voted for a value less than or equal to $\sigma^{\prime}$. Else $\sigma$ remains unchanged.
Thus, to change $\sigma$ from $\frac{2}{3}$ to $\frac{3}{4}$, the approval of a $\frac{3}{4}$-supermajority among $P$ is needed, and to change to unanimity ($\sigma=\frac{2|P|-1}{2|P|}$), approval by all members of $P$ is needed. And if the prevailing constitution states unanimity, then changing back to a $\frac{2}{3}$-supermajority requires unanimous consent among $P$.
The $h$-rule ensures that changing the supermajority threshold requires meeting that very threshold: To raise $\sigma$ to $\sigma^{\prime}$, you need a $\sigma^{\prime}$-supermajority supporting it (preventing minorities from imposing stricter requirements on the majority), while to lower $\sigma$ from the current value, you need the current $\sigma$-supermajority to agree (preventing minorities from weakening the constitution’s protection).
(3) Amending $\Delta$. The value of the parameter $\Delta$ should be increased or decreased only if the majority of the correct agents wish so. To this end, we offer the Suppress Outer-f parameter update rule [^1] In reference [32] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion. by Shahaf et al. [32], as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified. The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good” [29].
If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Else $\Delta$ remains unchanged.
We note several desirable properties of the Suppress Outer-f rule: If all Byzantines wish to increase $\Delta$ to slow-down the protocol, and the correct agents do not, then $\Delta$ will not be increased. Similarly, if all Byzantines wish to decrease $\Delta$ to undermine the liveness of the protocol, and the correct agents do not, then $\Delta$ will not be decreased. If all correct agents unanimously wish to update $\Delta$ to $\Delta^{\prime}$ then $\Delta$ will change to $\Delta^{\prime}$ no matter what the Byzantine agents vote, provided $f<\frac{1}{3}n$. In summary:
A constitution $(P,\sigma,\Delta)$ could be amended democratically to $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ as follows: 1. $P$ to $P^{\prime}$ via the approval of both a $\sigma$-supermajority among $P$ and all members of $P^{\prime}\setminus P$. 2. $\sigma$ to $\sigma^{\prime}$ via the h-rule for constitutional amendment [1]. 3. $\Delta$ to $\Delta^{\prime}$ via the suppress outer-f parameter update rule [32].
5.3 Constitutional Amendments
While we suggest above democratic decision-making processes for amending the three components of the constitution, their specifics are immaterial to the Constitutional Consensus protocol. The only thing that matters is the interface between the decision making process and the protocol, which takes the form of constitutional amendment decisions. The following definition assumes a Constitutional Consensus protocol instance identifier Id; we discuss its selection below.
Definition 5.1 (Constitutional Amendment Decision).
A constitutional amendment decision is a pair $d=(h,x)$, where $x=(\textsc{amend},\textit{Id},i,C,C^{\prime})$, Id is an instance identifier, $i\geq 1$ is the index of the decision within this instance, $C,C^{\prime}$ are constitutions, referred to as the old and new constitutions of $d$, respectively, and $h$ is a multisignature of the hash of $x$ by some $Q\subset\Pi$. If $i=1$ then $C=\bot$ and $P_{1}$ are referred to as the founders of the protocol instance Id—they are the signatories of its initial constitution and the participants of the first epoch.
Note that the structure of $d$ is that of a constitutional genesis block. During a run of a protocol instance Id, agents produce a possibly-infinite ordered sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$ using democratic decision-making processes discussed above or others, where $d_{i}=(h_{i},x_{i})$ and $x_{i}=(\textsc{amend},\textit{Id},i,C_{i-1},C_{i})$. The protocol commences with $d_{1}$ as the constitutional genesis block of the first epoch, and with each subsequent $d_{i}$ serving as the constitutional genesis block of the $i^{th}$ epoch.
Definition 5.2 (Valid Constitutional Amendment Decision and Block).
Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, the amendment $d_{i}$ with prevailing constitution $C_{i-1}=(P_{i-1},\sigma_{i-1},\Delta_{i-1})$, new constitution $C_{i}=(P_{i},\sigma_{i},\Delta_{i})$, and signatories $Q_{i}$ is valid if $i=1$, $P_{0}=\emptyset$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid and the signatories $Q_{i}$ are (i) a $\sigma_{i-1}$-supermajority in $P_{i-1}$, (ii) a $\sigma_{i}$-supermajority in $P_{i}$, and (iii) $(P_{i}\setminus P_{i-1})\subseteq Q_{i}$. The sequence is valid if all its members are valid. Two non-identical constitutional amendment decisions are conflicting if they have the same instance identifiers and the same index. Given a sequence of valid constitutional amendment decisions, a block is constitutional if it includes a valid constitutional amendment decision as payload.
Note that for agents to determine that a constitutional amendment decision is valid, they need access to all its predecessors. Also note that any two consecutive supermajorities may overlap, even be identical. In particular, if only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed. If only $\sigma_{i}$ changes then a single $\max(\sigma_{i-1},\sigma_{i})$-supermajority among $P_{i}$ is needed; this ensures that the $h$-rule [1] is followed. If only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$, a $\sigma_{i}$-supermajority among $P_{i}$, and all new members in $P_{i-1}$, are needed.
We assume that correct agents do not sign invalid or two conflicting constitutional amendments, and hence can make the following observation.
{observation}
[Safety of Constitutional Amendments] Given a valid finite sequence of constitutional amendment decisions $D=d_{1},\ldots d_{i}$, $i\geq 1$, if every $C_{i}$ in $D$ is safe, then the constitutional amendment process cannot extend $D$ with two conflicting valid constitutional amendments.
Proof 5.3.
By a counting argument, for this to happen a correct agent would have to sign two conflicting constitutional amendments, a contradiction.
Note that while the safety depends only on $\sigma_{i}$ and $P_{i}$, validity depends also on $\sigma_{i-1}$ and $P_{i-1}$, in order to adhere to the $h$-rule.
Grassroots selection of protocol instance identifier. Each constitutional amendment decision is associated with an instance Id of the Constitutional Consensus protocol. In a grassroots setup, the identifier of each instance should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents, and should be chosen without central authority, central resource, or coordination among disjoint instances. This can be achieved, for example, by having the identifier Id of a protocol instance consist of (the hash of) the founders (initial population) of the instance, each paired with the maximal serial number of any Constitutional Consensus protocol instance they have founded. This ensures Id to be unique if the founders include at least one correct agent, who would not sign the initial constitutional decision if its Id is not formed correctly.
Epochs and agent responsiveness. The standard assumption in consensus protocols is that all correct agents are responsive throughout the run. However, this assumption is unreasonable in a protocol like Constitutional Consensus in which the set of participating agents changes dynamically, especially since the set $\Pi$ of potential participants is infinite. Hence, we only assume that a correct agent $p$ is responsive only during an epoch in which it participates. The one exception is the case of agents that are behind, still in “older” epochs. To complete the epoch they may need to receive such blocks from former correct participants in that epoch. Hence, agent may still need to respond to nack-blocks pertaining to blocks they have issued in past epoch.
Still, an agent may determine that en epoch is ‘closed’ (and no further nack-blocks need to be served from it) if it has evidence that all participants in the epoch have ended it. In particular, an epoch can be ‘closed’ when all its participants have coronated the constitutional amendment decision that ended (see End Epoch below).
Note that the interaction in which a non-responsive agent $p$ needs to be ‘woken up’ in order to sign a constitutional amendment in which $p$ is a new participant happens outside the Constitutional Consensus protocol. In the practical settings we envision this would happen through social interactions (in-person or digital) between existing participants and prospective new participants.
5.4 Epoch Transitions via Constitutional Amendments
Protocol concepts. Every epoch commences with a constitutional genesis block that specifies its prevailing constitution. If, during the epoch, members of the prevailing population become aware of a valid decision $d$ to amend the prevailing constitution to a new constitution, they cease normal operation and only endeavour to finalize this decision. Once an agent has a wave $W$ that finalizes a constitutional block with $d$, it coronates $d$ by producing a coronate block with payload $(\textsc{coronate},d)$, sending it to both the old and new populations of $d$, and ceasing participation in the epoch. Members of the old population that receive it and have yet to coronate $d$ can nack any missing blocks to obtain $W$ and coronate $d$ themselves. A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block. An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own (implying that $d$ is final in its blocklace).
The protocol. The protocol produces blocks also with constitutional amendment decisions as payload; such a block is referred to as a constitutional block. When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution. Subsequent ordered blocks with the same amendment are suppressed from the output.
The protocol is extended in Algorithm 2 as follows:
  1. 1.
    Amendment Finalization occurs whenever the agent $p$ knows of a valid constitutional amendment $d$ that has not been finalized, upon which it continues following the protocol but only producing constitutional blocks with $d$ as a payload, until $d$ is finalized. Here, knows has a dual meaning: Either $p$ knows that the external democratic decision making process resulted in a new valid constitutional amendment $d$, or $p$ has a constitutional block with $d$. In either case, no constitutional block with $d$ has been finalized yet in $B$.
  2. 2.
    End Epoch applies when a constitutional block with $d$ is finalised, upon which members of the old population create a coronate block with $d$, send it to all members of the old and new populations, and cease participation in the old epoch. Coronation also plays a role in the democratic decision process: A correct agent would not sign a new constitutional amendment decision before coronating the last constitutional amendment it was a signatory to.
  3. 3.
    Start Epoch applies when any agent $p$ in the new population receives from the old population a supermajority of coronate block with decision $d$ (and provided $p$ issued a coronate block if it is also a member of the old population), upon which $p$ creates a constitutional genesis block $b$ with $d$ and updates its epoch identifier to be $b$.
  4. 4.
    Respond to Belated nack applies when any agent $p$ receives a nack $q$-block $b$ for a $p$-block of a past epoch $e$, upon which $p$ responds to the nack to help $q$ complete this past epoch. Note that if $p$ had previously received a coronate $q$-block ending $e$, then $p$ knows that $q$ is not missing any epoch $e$ blocks; in this case “judicially send…” would be a no-op. Moreover, an agent may ‘write off’ an epoch $e$ as complete after receiving from all participants in $e$ blocks coronating the end of $e$.

### v5

@@ E.2. Liveness
Observation E.9 (Single-epoch Liveness).
A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
Theorem E.10.
@@
Proof.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $b_{p}]$) have not been received by $q$, then after receiving $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction. ∎
@@
Observation E.12 (Sequential Advance).
If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
@@
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation E.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition E.11, this means that no correct agent sends blocks in rounds >r>r.
Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition E.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
By Proposition E.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
@@
Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction. ∎
@@
Proposition E.14 (Finite implies quiescence).
@@
Proposition E.15 (Block issuance).
@@
Proof.
Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload. First, consider the case that $p$’s rounds do not advance in some suffix of the run. Then by Proposition E.14, $p$’s blocklace is quiescent. Hence, condition 1 of the Issue rule is satisfied, and the payload is sent. A contradiction. Therefore, $p$’s advanced rounds increase indefinitely. Consider a time when $p$’s highest advanced round $r>r_{0}$ and the payload has not been sent. Then $p$ issues a block either in round $r+1$ (by the Issue rule) or in round $r$ (by the Issue backlog rule). A contradiction. ∎
Proposition E.16 (Transaction Liveness in finite runs).
@@
Proof.
By Proposition E.11, $B$ includes all ordinary blocks sent by correct agents. By Proposition E.14, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$. Moreover, by Proposition E.15, no agents have payloads to send. It follows that all input transactions have been sent in ordinary blocks and ordered. ∎
Proposition E.17 (Epoch Liveness in finite runs).
@@
Proof.
Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run. Because $t$ is at some point final in $p_{1}$’s blocklace, its blocklace eventually includes a block $b$ with $t$ in its payload and a super-majority of blocks that observe $b$. This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$. By Proposition E.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$. Because $B_{2}$ is finite in a suffix of the run, by Proposition E.14, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed. ∎
@@
Proposition E.18 (First round progress after GST).
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt¿GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof.
@@
When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block. ∎
@@
Proposition E.19 (Progress after GST).
@@
Proof.
By Proposition E.18, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$. This means that it is impossible to create a valid round $r+1$ block that does not endorse the leader, as such a block would have to point to a super-majority of round $r$ blocks excluding the leader block. Therefore, every valid round $r+2$ block ratifies the leader block, and the block is finalized in round $r+2$ as soon as it advances at any correct agent. ∎
@@
Proposition E.20 (Infinite quiescent waves).
@@
Proof.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition E.11, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalising, and finalization is monotonic, there are infinitely many finalising waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
@@
Consider a finalising wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition E.11). If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
@@
So when the third round of wave $k+1$ advances, every correct agent either observes a super-majority of blocks that ratify the leader block or observes a block whose closure is quiescent. In the former case the leader block is finalized, ordering all preceding ordinary blocks including $b_{2}$, and in the latter $b_{2}$ is already final in the block’s closure. ∎
@@
By Observation E.9, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent. By Proposition E.15, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks. Similarly, every output transaction of a correct agent is included in some ordinary block. Let $b$ be a non-empty ordinary $p$-block included in some correct agent’s blocklace in a run of Constitutional Consensus. If $p$ is correct, then by Proposition E.11, every correct agent receives and incorporates $b$ in its local blocklace. Otherwise, $b$ is finalized in some blocklace, implying that it is approved by at least one $q$-block $b^{\prime}$ by a correct agent $q$. Again, by Proposition E.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
@@
We consider three cases: First, assume the run is finite. By Propositions E.16 and E.17, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds. Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. By Proposition E.20, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
@@
Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time. Let $k$ be a wave whose last round advances after GST and after which no wave is quiescent at any correct agent’s blocklace. By fairness of leader selection, there are inifinitely many waves $k^{\prime}>k$ with a correct leader. By Proposition E.19, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds. ∎
Appendix F The Amendment Setting
We describe the general democratic operation.
Unified proposing and voting. Following Bulteau et al. (Bulteau et al., 2021), there is no proposal phase. Whenever an agent decides to vote (afresh or as update), it issues a block with its timestamped and signed vote. A vote is persistent: if agent $i$ does not update its vote, the most recent one in the blocklace remains in effect. The prevailing constitution thus always reflects the current preferences of all agents without requiring repeated re-voting. As we show next, the decision spaces of $P$, $\sigma$, and $\Delta$ are all naturally embedded in corresponding metric spaces.
Domain structure of the parameters. The three constitutional parameters have distinct domain structures, which determine the appropriate aggregation rule for each:
  • $P$ (population): For each potential participant $p$, each agent votes yes or no—effectively, a separate referendum on each member. This is the simplest form of status-quo voting.
  • $\sigma$ (supermajority threshold): Agents state a preferred value in $[0.5,1)$, a naturally single-peaked domain. Additionally, $\sigma$ is self-referential: it is the threshold used to amend the constitution, including $\sigma$ itself. The amendment rule must therefore be consistent with its own output.
  • $\Delta$ (timeout): Agents state a preferred value in $\mathbb{R}^{+}$, again a single-peaked one-dimensional domain.
Appendix G Constitutional Amendment Pseudocode